# Text-Aware Diffusion for Policy Learning

### v1

@@ 5 Conclusion and Future Work
We present Text-Aware Diffusion for Policy Learning (TADPoLe), a framework that optimizes a policy according to a provided natural language prompt through a pretrained text-conditioned diffusion model. TADPoLe enables novel behaviors to be learned in a zero-shot manner purely from text conditioning, and also offers a promising angle to train policies to behave in accordance with natural priors summarized from large-scale pretraining. TADPoLe can be applied across visual environments and different robotic states without modification, and we experimentally demonstrate that TADPoLe is able to learn to novel goal-achievement as well as continuous locomotion behaviors conditioned only on text, across Humanoid, Dog, and Meta-World environments.
Limitations: An observed limitation of TADPoLe is that it is difficult to explicitly control the weight each individual word of an input text prompt has on the reward provided to the agent. For certain prompts, TADPoLe could potentially cause the agent to remain stationary since it may focus on alignment with the noun in the phrase rather than details of the goal. How to provide fine-grained control over the text-conditioning is an interesting direction to explore in future work. Further interesting future work include utilizing multiple camera views simultaneously to compute the dense reward, as environments generally allow flexible rendering from arbitrary angles. Another observed limitation is that TADPoLe depends on a highly stochastic operation, namely, repeatedly resampling a Gaussian source noise vector at each timestep. The behavior of the resulting policy, after training for many iterations, can therefore vary for the same input text prompt, and potentially cause high variance in both visual and quantitative performance. How to control the stability of convergence to a consistent policy across repeated runs is an interesting future direction for exploration.
Acknowledgements. We would like to thank Amil Merchant, Daniel Ritchie, David Bau, Ben Poole, Ting Chen, Nate Gillman, and Yilun Du for helpful discussions and feedback. This work is supported by Samsung Advanced Institute of Technology and a Richard B. Salomon Faculty Research Award for Chen Sun. Our research was conducted using computational resources at the Center for Computation and Visualization at Brown University.
@@ Appendix A Intuition Regarding A Reasonable Noise Level Range
For a sufficiently low noise value, the conditional and unconditional noise prediction is similar for a particular text prompt and rendered frame that already has high cross-modal alignment. Being already close in data space to the desired mode described by the text prompt, the unconditional prediction will also cheaply seek to denoise the input towards the original input. Therefore, too low a noise value may cause the computed reward signal to decrease as cross-modal alignment increases, which runs counter to what is desirable. This is supported by looking at the left tails of the two graphs depicted in Figure A1, where the difference in computed TADPoLe reward between well-aligned paired inputs and misaligned paired inputs is small.
Similarly, choosing a sufficiently high noise value may also cause unconditional and conditional predictions to be similar. For a given noisy input with virtually all of the spatial structure perturbed, the pretrained denoising model intuitively makes denoising predictions that fill in general structure to the image irrespective of text-conditioning. This is intuitively similar to the behavior of an unconditional prediction. Once again, we can visually observe this in the right tails of both graphs in Figure A1; for high noise values, both misaligned and well-aligned paired inputs have similar computed TADPoLe reward.
@@
Figure A1: Noise range intuition for a fixed image but two distinct prompts (left), and for a fixed prompt but two distinct images (right). Through visualization, we verify that $U(400,500)$ is a reasonable range from which to sample noise levels that can meaningfully distinguish text-alignment for arbitrarily rendered frames.
@@ Appendix C Meta-World Tasks
Success Rate (%) Door Open Door Close Drawer Open Drawer Close Window Open Window Close Coffee Push Button Press Soccer Peg Insert Side Shelf Place Lever Pull Average
Sparse 0 ($\pm$ 0) 99.7 ($\pm$ 1.0) 74.7 ($\pm$ 28.7) 99.3 ($\pm$ 2.0) 94.3 ($\pm$ 5.8) 93.0 ($\pm$ 16.9) 46.3 ($\pm$ 11.3) 86.7 ($\pm$ 29.1) 12.7 ($\pm$ 14.1) 0 ($\pm$ 0) 0 ($\pm$ 0) 0 ($\pm$ 0) 50.6
VLM-RM 0 ($\pm$ 0) 79.7 ($\pm$ 39.8) 10.0 ($\pm$ 30.0) 100 ($\pm$ 0) 9.7 (±$\pm 29.0$) 0 ($\pm$ 0) 4.0 ($\pm$ 12.0) 30.0 ($\pm$ 45.8) 5.3 ($\pm$ 11.6) 0 ($\pm$ 0) 0 ($\pm$ 0) 11.3 ($\pm$ 18.3) 20.8
TADPoLe 40.0 ($\pm$ 49.0) 100 ($\pm$ 0) 45.3 ($\pm$ 46.6) 100 ($\pm$ 0) 74.0 ($\pm$ 37.9) 30.0 ($\pm$ 45.8) 18.6 ($\pm$ 20.5) 73.0 ($\pm$ 38.5) 25.0 ($\pm$ 15.4) 0 ($\pm$ 0) 0 ($\pm$ 0) 0 ($\pm$ 0) 42.2

### v2

@@ 5 Conclusion and Future Work
We present Text-Aware Diffusion for Policy Learning (TADPoLe), a framework that optimizes a policy according to a provided natural language prompt through a pretrained text-conditioned diffusion model. TADPoLe enables novel behaviors to be learned in a zero-shot manner purely from text conditioning, and also offers a promising angle to train policies to behave in accordance with natural priors summarized from large-scale pretraining. TADPoLe can be applied across visual environments and different robotic states without modification, and we experimentally demonstrate that TADPoLe is able to learn novel goal-achievement as well as continuous locomotion behaviors conditioned only on text, across Humanoid, Dog, and Meta-World environments.
Limitations: An observed limitation of TADPoLe is that it is difficult to explicitly control the weight each individual word of an input text prompt has on the reward provided to the agent. For certain prompts, TADPoLe could potentially cause the agent to remain stationary since it may focus on alignment with the noun in the phrase rather than details of the goal. How to provide fine-grained control over the text-conditioning is an interesting direction to explore in future work. Further interesting future work includes utilizing multiple camera views simultaneously to compute the dense reward, as environments generally allow flexible rendering from arbitrary angles. Another observed limitation is that TADPoLe depends on a highly stochastic operation, namely, repeatedly resampling a Gaussian source noise vector at each timestep. The behavior of the resulting policy, after training for many iterations, can therefore vary for the same input text prompt, and potentially cause high variance in both visual and quantitative performance. How to control the stability of convergence to a consistent policy across repeated runs is an interesting future direction for exploration.
Acknowledgements. We would like to thank Amil Merchant, Daniel Ritchie, David Bau, Ben Poole, Ting Chen, Nate Gillman, and Yilun Du for helpful discussions and feedback. This work is supported by Samsung Advanced Institute of Technology, NASA, and a Richard B. Salomon Faculty Research Award for Chen Sun. Our research was conducted using computational resources at the Center for Computation and Visualization at Brown University.
@@ Appendix A Intuition Regarding A Reasonable Noise Level Range
For a sufficiently low noise value, the conditional and unconditional noise prediction is similar for a particular text prompt and rendered frame that already has high cross-modal alignment. Being already close in data space to the desired mode described by the text prompt, the unconditional prediction will also cheaply seek to denoise the input towards the original input. Therefore, too low a noise value may cause the computed reward signal to decrease as cross-modal alignment increases, which runs counter to what is desirable. This is supported by looking at the left tails of the two graphs depicted in Figure A1, where the difference in computed TADPoLe rewards between well-aligned paired inputs and misaligned paired inputs is small.
Similarly, choosing a sufficiently high noise value may also cause unconditional and conditional predictions to be similar. For a given noisy input with virtually all of the spatial structure perturbed, the pretrained denoising model intuitively makes denoising predictions that fill in general structure to the image irrespective of text-conditioning. This is intuitively similar to the behavior of an unconditional prediction. Once again, we can visually observe this in the right tails of both graphs in Figure A1; for high noise values, both misaligned and well-aligned paired inputs have similar computed TADPoLe rewards.
@@
Figure A1: Noise range intuition for a fixed image but two distinct prompts (left), and for a fixed prompt but two distinct images (right). Through visualization, we verify that $U(400,500)$ is a reasonable range from which to sample noise levels that can meaningfully distinguish vision-text alignment for arbitrarily rendered frames.
@@ Appendix C Meta-World Tasks
Success Rate (%) Door Open Door Close Drawer Open Drawer Close Window Open Window Close Coffee Push
Sparse 0 ($\pm$ 0) 99.7 ($\pm$ 1.0) 74.7 ($\pm$ 28.7) 99.3 ($\pm$ 2.0) 94.3 ($\pm$ 5.8) 93.0 ($\pm$ 16.9) 46.3 ($\pm$ 11.3)
VLM-RM 0 ($\pm$ 0) 79.7 ($\pm$ 39.8) 10.0 ($\pm$ 30.0) 100 ($\pm$ 0) 9.7 (±$\pm 29.0$) 0 ($\pm$ 0) 4.0 ($\pm$ 12.0)
TADPoLe 40.0 ($\pm$ 49.0) 100 ($\pm$ 0) 45.3 ($\pm$ 46.6) 100 ($\pm$ 0) 74.0 ($\pm$ 37.9) 30.0 ($\pm$ 45.8) 18.6 ($\pm$ 20.5)
Success Rate (%) Button Press Soccer Peg Insert Side Shelf Place Lever Pull Average
Sparse 86.7 ($\pm$ 29.1) 12.7 ($\pm$ 14.1) 0 ($\pm$ 0) 0 ($\pm$ 0) 0 ($\pm$ 0) 50.6
VLM-RM 30.0 ($\pm$ 45.8) 5.3 ($\pm$ 11.6) 0 ($\pm$ 0) 0 ($\pm$ 0) 11.3 ($\pm$ 18.3) 20.8
TADPoLe 73.0 ($\pm$ 38.5) 25.0 ($\pm$ 15.4) 0 ($\pm$ 0) 0 ($\pm$ 0) 0 ($\pm$ 0) 42.2
@@ Appendix D Training Curves
Table A7: Numerical results from the user study on text-alignment. Out of 25 anonymous users, for each method, we report the percentage that believe the video achieved by a policy at the end of training is aligned with the provided text prompt. Entries above 50% are displayed in Table 1 as checkmarks (); below 50% are displayed as x-marks ().
Environment Prompt VLM-RM LIV Text2Reward TADPoLe (Ours)
Humanoid “a person in lotus position" 52% 28% 4% 60%
Humanoid “a person doing splits" 64% 88% 0% 84%
Humanoid “a person kneeling" 48% 4% 0% 64%
Dog “a dog standing" 16% 0% 40% 84%